Zhao Co. has fixed costs of $403,200. Its single product sells for $183 per unit, and variable costs are $120 per unit. If the company expects sales of 10,000 units, compute its margin of safety in dollars and as a percent of expected sales.
Answer #1
Selling price per unit 183
Less: Variable cost per unit 120
Contribution margin per unit 63
Fixed costs 403200
/Contribution margin per unit 63
Break even point in units 6400
Margin of safety in units 3600 =10000-6400
Margin of safety in dollars 658800 =3600*183
Margin of safety as a percentage of sales 36% =3600/10000